# Anomalous radiation reaction in a circularly polarized field

### v6

@@ Abstract
Introduction. The laws of nature describing the motion of charged particles in electromagnetic fields lie in the basis of most physical processes in the Universe and, therefore, they are permanently in the focus of attention despite their long prehistory. Since a charged particle driven by a high-frequency field makes oscillating movement, it emits electromagnetic radiation. If the field is strong enough, the radiation reaction dominates the particle motion and needs the quantum electrodynamics (QED) to be described accurately. Although this research area is exciting nowadays due to achievements in the laser generation of strong fields Gonoskov_2022 ; Fedotov_2023 ; Piazza_2012 ; Piazza_2008 ; Piazza_2010_1 ; Piazza_2010 ; Piazza_2009 ; Ilderton_2013 ; Cole_2018 ; Poder_2018 , the features of the QED corrections depending on the field polarization still wait for detailed analysis. To fill partially this gap, the emission of photons by an electron rotating under a circularly polarized field was considered within the Floquet theory which is conventionally used to describe various periodically driven quantum systems Goldman_2014 ; Bukov_2015 ; Eckardt_2015 ; Casas_2001 ; Kibis_2022 . Solving this problem, it was found that the photon emission leads to the quantum recoil force acting on the emitting electron perpendicularly to the velocity of its forward movement. This result is surprising since the classical Lorentz-Abraham-Dirac (LAD) force — which was introduced into physics at the dawn of electrodynamics to describe the radiation reaction and takes deserved place in textbooks Landau_2 ; Barut_book — is directed oppositely to the velocity. The present Letter is devoted to the first theoretical analysis of this anomalous radiation reaction which can manifest itself in strong laser fields.
Model. For definiteness, let us consider a classically strong circularly polarized homogeneous electromagnetic field with the zero scalar potential and the vector potential
@@
where $E_{0}$ is the electric field amplitude, $\omega$ is the field frequency, $c$ is the light speed in vacuum, and the field polarization is assumed to be clockwise. Within the framework of classical electrodynamics, the vector potential (A1) describes the rotating electric field, $\mathbf{E}=-(1/c)\partial_{t}\mathbf{A}$, which induces electron rotation with the velocity $v_{0}=eE_{0}/m_{e}\omega$ along circular trajectory of the radius $r_{0}=|v_{0}|/\omega$, where $m_{e}$ is the electron mass and $e$ is the electron charge Landau_2 . Proper generalization of the model field (A1) to the physically important case of electromagnetic wave will be done in the following. The non-relativistic quantum dynamics of an electron under the field (A1) is defined by the Hamiltonian $\hat{\cal H}_{0}=\left(\hat{\mathbf{p}}-{e}\mathbf{A}/c\right)^{2}/{2m_{e}},$ where $\hat{\mathbf{p}}=(\hat{{p}}_{x},\,\hat{{p}}_{y},\,\hat{{p}}_{z})$ is the momentum operator. The exact eigenfunctions of the Hamiltonian are
@@
The interaction between an electron and a photon in the presence of the background field (A1) is described by the Hamiltonian $\hat{\cal H}=\hat{\cal H}_{0}+\hat{V}^{(\pm\mathbf{q})}$, where $\hat{V}^{(\pm\mathbf{q})}$ is the Hamiltonian of electron-photon interaction, and the signs “++” and “-” correspond to the absorption and emission of a photon with the wave vector $\mathbf{q}$, respectively. Since the Floquet functions (A3) form the complete orthonormal function system for any time $t$, the solution of the Schrödinger problem with the Hamiltonian $\hat{\cal H}$ can be sought as an expansion $\psi=\sum_{\mathbf{k}^{\prime}}c_{\mathbf{k}^{\prime}}(t)\psi_{\mathbf{k}^{\prime}}.$ Substituting this expansion into the Schrödinger equation, $\hat{\cal H}\psi=i\hbar\partial_{t}\psi$, we arrive at the Schrödinger equation written in the Floquet representation,
@@
Results and discussion. Using the Floquet functions (A3) as a basis to expand the electron wave function, we take into account the electron interaction with the strong background field (A1) accurately (non-perturbatively), whereas the electron interaction with photons can be considered as a weak perturbation. First of all, let us consider the one-vertex photon emission pictured in Fig. 1a, assuming an electron to be initially in the Floquet state $\psi_{\mathbf{k}}$. Solving the Schrödinger equation (A6) within the first order of the conventional perturbation theory Landau_3 , we arrive at the amplitude of the one-vertex process,
@@
Applying Eq. (A14), one can find the total power that an electron radiates,
@@
where $\gamma=1/\sqrt{1-{v}^{2}/c^{2}}$ is the Lorentz factor. Indeed, substituting the velocity (A4) into Eq. (A23) and averaging it over the field period, $2\pi/\omega$, we arrive at the same Eq. (A22). As a consequence, Eq. (A22) exactly coincides also with the Landau-Lifshitz equation describing the radiation reaction in an electromagnetic field Landau_2 . To extend the developed theory for an electron in an electromagnetic wave, one needs to make the replacement $\omega t\rightarrow\omega t-\mathbf{q}_{0}\mathbf{r}$ in the vector potential (A1), where $\mathbf{q}_{0}=(0,\,0,\,\omega/c)$ is the wavevector of the wave. It can be easily proved that the Floquet functions (A3) with the same replacement stay eigenfunctions of the Hamiltonian $\hat{\cal H}_{0}$ in the non-relativistic limit ($v/c,\,\hbar\omega/m_{e}c^{2}\ll 1$). As an only consequence of such a modification of them, the replacement of the delta function, $\delta^{3}(\mathbf{k}-\mathbf{k}^{\prime}\pm\mathbf{q})\rightarrow\delta^{3}(\mathbf{k}-\mathbf{k}^{\prime}\pm\mathbf{q}+m\mathbf{q}_{0})$, should be done in the matrix element (A9). Physically, this means that emission of each photon at the main radiation harmonic with $m=1$ is accompanied by transfer of the momentum $\hbar\mathbf{q}_{0}$ from the wave to the electron (the Compton scattering). Therefore, the photon drag force,
@@
It follows from the aforesaid that the one-vertex emission process pictured in Fig. 1a leads to exactly the same physical results as the classical electrodynamics. To find the first quantum correction to them, one needs to take into account the one-loop process pictured in Fig. 1b. Solving the Schrödinger equation (A6) within the third order of the conventional perturbation theory Landau_3 , we arrive at the amplitude of the one-loop process,
@@
is the QED correction to the probability (A14) arisen from the one-loop process SM . Applying Eq. (A26), the total power emitted by an electron can be written as
@@
It should be noted that the emission processes described by the loop and loopless Feynman diagrams make substantially different contributions to the radiation reaction (A35). Mathematically, this follows from the fact that the anomalous recoil force (A36) appears due to the singularities of the probability amplitude (A27), which arise from virtual photons SM . In contrast to the loop diagrams, the loopless ones do not involve virtual photons and, therefore, their amplitudes are devoid of the singularities. As a consequence, the two emission processes pictured in Fig. 1 describe the radiation reaction (A35) accurately within the leading order of electron-photon interaction: The one-vertex process makes main contribution to the normal recoil force (A22), whereas the one-loop process is responsible for the anomalous force (A35). It should be stressed that the one-vertex photon emission pictured in Fig. 1a is the direct emission process which does not involve intermediate electron-photon states and, therefore, is similar physically to the classical process of electromagnetic emission. As a consequence, it results in the classical recoil force (A22), where the Planck constant, $\hbar$, is cancelled. As to the one-loop photon emission pictured in Fig. 1b, it is the indirect process going through the intermediate state with the virtual photon $\mathbf{q}^{\prime}$. It has no analogue within classical electrodynamics and, therefore, yields the anomalous recoil force (A36) involving the fine structure constant, $\alpha=e^{2}/\hbar c$, which is of purely quantum nature.
@@
Evolution of the electron velocity $\mathbf{v_{k}}$ under the force (A35) can be described by the Newton equation, $m_{e}\dot{\mathbf{v}}_{\mathbf{k}}=\mathbf{F}_{\|}+\mathbf{F}_{\perp}$, which leads to the electron trajectory pictured schematically in Fig. 3. Since the quantum recoil force, $\mathbf{F}_{\perp}$, is directed perpendicularly to the electron velocity, it does not decelerate an electron in contrast to the classical recoil force, $\mathbf{F}_{\|}$, but bends its trajectory similarly to the Lorentz force acting on an electron in a magnetic field. Such an analogy has the deep physical nature since both a high-frequency circularly polarized field and a stationary magnetic field break the time-reversal symmetry. As a consequence, a circularly polarized field can induce various magnetic-like effects, including the spin polarization Kibis_2022 , bandgap opening in solids Kibis_2010 ; Wang_2013 ; Sie_2015 ; Cavalleri_2020 , persistent currents in conductors Kibis_2011 ; Melnikov_2021 , etc. It should be stressed that the spatial asymmetry of photon emission responsible for the quantum recoil force (A36) — which is described by the last term in Eq. (III) — also arises from the broken time-reversal symmetry. Therefore, the discussed anomalous radiation reaction acting on a charged particle perpendicularly to its velocity takes place in any electromagnetic field with nonzero angular momentum $\mathbf{L}$ (circularly and elliptically polarized fields) and vanishes only in the case of linearly polarized field with $\mathbf{L}=0$. It follows from the aforesaid, particularly, that Eq. (A36) should be considered as a quantum correction to the Landau-Lifshitz equation Landau_2 describing the classical radiation reaction force in an electromagnetic field of most general form. It should be noted also that the angular momentum of a circularly polarized field, $\mathbf{L}$, is codirectional to the angular momentum of a charge rotating under the field. This means that the quantum recoil force (A36) is similar phenomenologically to the classical Magnus force, $\mathbf{F}_{M}\propto[\mathbf{L}\times\mathbf{v}]$, acting on a body which rotates with the angular momentum $\mathbf{L}$ and makes forward movement with the velocity $\mathbf{v}$ in a medium with friction (see, e.g., Ref. Clancy_book, ). Treating the classical recoil force (A22) as a radiation friction, the quantum recoil force (A36) can be considered qualitatively as a specific kind of the Magnus force arisen from this friction acting on a rotating charge.
@@
Although the present theory is developed for the homogeneous field (A1), the force (A36) very weakly depends on spatial profile of the field if its characteristic scale much exceeds the radius of circular electron trajectory, $r_{0}=|v_{0}|/\omega$. As to the time profile of the field (the laser pulse duration), its characteristic scale should much exceed the radiative life time (A20). Since the radiative life time, $\tau$, must be large enough to satisfy the condition $\omega\tau\gg 1$, it is most suitable to detect the force (A36) with non-relativistic laser fields which can be easily generated in the continuous-wave regime. Discussing dependence of the force (A36) on the field strength, $E_{0}$, it should be noted that the Floquet functions (A3) are exact solutions of the non-relativistic Schrödinger problem for an electron in the circularly polarized background field (A1) of any strength. Therefore, the developed theory is applicable to the background field of the strength $E_{0}$ satisfying only the non-relativistic condition, $v_{0}/c=eE_{0}/m_{e}\omega c\ll 1$. To keep the ratio $v_{0}/c$ to be small enough within applicability of the developed non-relativistic theory, one can consider laser fields of the strength $E_{0}$ up to $10^{11}$ V/m. Estimating the force (A36) quantitatively, it should be noted that the electric field amplitude $E_{0}$ can reach $10^{10}$ V/m for an electromagnetic wave generated by an ordinary hundred-kilowatt laser with the wavelengths $\lambda\sim 10^{-6}$ m. In such a field, the quantum force (A36) yields the electron acceleration $a_{\perp}=F_{\perp}/m_{e}\sim 10^{3}(v_{\mathbf{k}}/c)$ m/s2, which can be macroscopically large even for non-relativistic electrons satisfying the condition $v_{\mathbf{k}}\ll c$. Thus, the discussed anomalous radiation reaction is the rare case of macroscopic QED effect which can be observable for slow electrons in laser fields of non-relativistic intensity. As to analysis of fast electrons moving in ultra-strong fields, it needs the relativistic extension of the developed Floquet theory based on the Dirac equation SM .
@@
Conclusion. It is shown theoretically that the emission of photons by an electron rotating under a high-frequency circularly polarized electromagnetic field yields the quantum recoil force acting on the electron perpendicularly to the velocity of its forward movement. In contrast to the classical Lorentz-Abraham-Dirac recoil force directed oppositely to the velocity, it does not decelerate the electron but substantially modifies its trajectory. Such an anomalous radiation reaction arises from the one-loop QED correction to the photon emission and can manifest itself in strong laser fields.
@@
Acknowledgments. This work was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” and the Ministry of Science and Higher Education of the Russian Federation (Project FSUN-2023-0006).

### v7

@@ Abstract
Introduction—The laws of nature describing the motion of charged particles in electromagnetic fields lie in the basis of most physical processes in the Universe and, therefore, they are permanently in the focus of attention despite their long prehistory. Since a charged particle driven by a high-frequency field makes oscillating movement, it emits electromagnetic radiation. If the field is strong enough, the radiation reaction dominates the particle motion and needs the quantum electrodynamics (QED) to be described accurately. Although this research area is exciting nowadays due to achievements in the laser generation of strong fields Gonoskov_2022 ; Fedotov_2023 ; Piazza_2012 ; Piazza_2008 ; Piazza_2010_1 ; Piazza_2010 ; Piazza_2009 ; Ilderton_2013 ; Cole_2018 ; Poder_2018 , the features of the QED corrections depending on the field polarization still wait for detailed analysis. To fill partially this gap, the emission of photons by an electron rotating under a circularly polarized field was considered within the Floquet theory which is conventionally used to describe various periodically driven quantum systems Goldman_2014 ; Bukov_2015 ; Eckardt_2015 ; Casas_2001 ; Kibis_2022 . Solving this problem, it was found that the photon emission leads to the quantum recoil force acting on the emitting electron perpendicularly to the velocity of its forward movement. This result is surprising since the classical Lorentz-Abraham-Dirac (LAD) force — which was introduced into physics at the dawn of electrodynamics to describe the radiation reaction and takes deserved place in textbooks Landau_2 ; Barut_book — is directed oppositely to the velocity. The present Letter is devoted to the first theoretical analysis of this anomalous radiation reaction which can manifest itself in strong laser fields.
Model—For definiteness, let us consider a classically strong circularly polarized homogeneous electromagnetic field with the zero scalar potential and the vector potential
@@
where $E_{0}$ is the electric field amplitude, $\omega$ is the field frequency, $c$ is the light speed in vacuum, and the field polarization is assumed to be clockwise. Within the framework of classical electrodynamics, the vector potential (1) describes the rotating electric field, $\mathbf{E}=-(1/c)\partial_{t}\mathbf{A}$, which induces electron rotation with the velocity $v_{0}=eE_{0}/m_{e}\omega$ along circular trajectory of the radius $r_{0}=|v_{0}|/\omega$, where $m_{e}$ is the electron mass and $e$ is the electron charge Landau_2 . Proper generalization of the model field (1) to the physically important case of electromagnetic wave will be done in the following. The non-relativistic quantum dynamics of an electron under the field (1) is defined by the Hamiltonian $\hat{\cal H}_{0}=\left(\hat{\mathbf{p}}-{e}\mathbf{A}/c\right)^{2}/{2m_{e}},$ where $\hat{\mathbf{p}}=(\hat{{p}}_{x},\,\hat{{p}}_{y},\,\hat{{p}}_{z})$ is the momentum operator. The exact eigenfunctions of the Hamiltonian are
@@
where $\mathbf{r}=(x,y,z)=(r\sin\theta\cos\varphi,\,r\sin\theta\sin\varphi,\,r\cos\theta)$ is the electron radius vector, $\theta$ and $\varphi$ are the polar and azimuth angle, respectively, $\mathbf{k}=(k_{x},\,k_{y},\,k_{z})=(k\sin\theta_{\mathbf{k}}\cos\varphi_{\mathbf{k}},\,k\sin\theta_{\mathbf{k}}\sin\varphi_{\mathbf{k}},\,k\cos\theta_{\mathbf{k}})$ is the electron wave vector, $\varepsilon_{\mathbf{k}}=\hbar^{2}{\mathbf{k}}^{2}/2m_{e}$ is the kinetic energy of forward movement of the electron, $\varepsilon_{0}=m_{e}v_{0}^{2}/2$ is the kinetic energy of electron rotation under the field (1), and $\cal{V}$ is the normalization volume. The wave functions (2) can be easily proved by direct substitution into the Schrödinger equation, $\hat{\cal H}_{0}\psi_{\mathbf{k}}=i\hbar\partial_{t}\psi_{\mathbf{k}}$, and should be treated as the Floquet functions Goldman_2014 ; Bukov_2015 ; Eckardt_2015 ; Casas_2001 describing dynamics of an electron strongly coupled to the field (1) (dressed by the field). The averaged electron velocity in the Floquet state (2) reads
@@
where $\mathbf{v_{k}}=\hbar\mathbf{k}/m_{e}$ and $\mathbf{v}_{0}=v_{0}(-\cos\omega t,\,-\sin\omega t,\,0)$ are the velocity of forward and rotational movement of the electron, respectively. As expected, the velocity (3) exactly coincides with the classical velocity of an electron under a circularly polarized field Landau_2 . Next, let us find corrections to this velocity arisen from the emission of photons by the dressed electron.
The interaction between an electron and a photon in the presence of the background field (1) is described by the Hamiltonian $\hat{\cal H}=\hat{\cal H}_{0}+\hat{V}^{(\pm\mathbf{q})}$, where $\hat{V}^{(\pm\mathbf{q})}$ is the Hamiltonian of electron-photon interaction, and the signs “++” and “-” correspond to the absorption and emission of a photon with the wave vector $\mathbf{q}$, respectively. Since the Floquet functions (2) form the complete orthonormal function system for any time $t$, the solution of the Schrödinger problem with the Hamiltonian $\hat{\cal H}$ can be sought as an expansion $\psi=\sum_{\mathbf{k}^{\prime}}c_{\mathbf{k}^{\prime}}(t)\psi_{\mathbf{k}^{\prime}}.$ Substituting this expansion into the Schrödinger equation, $\hat{\cal H}\psi=i\hbar\partial_{t}\psi$, we arrive at the Schrödinger equation written in the Floquet representation,
@@
is the transition current, and $\mathbf{j}_{\mathbf{k}^{\prime}\mathbf{k}}^{(m)}$ are its Fourier harmonics. Applying the Jacobi-Anger expansion, $e^{i\xi\sin\omega t}=\sum_{m=-\infty}^{\infty}J_{m}(\xi)e^{im\omega t}$, to find these harmonics, the matrix element (5) can be rewritten as
@@
Results and discussion—Using the Floquet functions (2) as a basis to expand the electron wave function, we take into account the electron interaction with the strong background field (1) accurately (non-perturbatively), whereas the electron interaction with photons can be considered as a weak perturbation. First of all, let us consider the one-vertex photon emission pictured in Fig. 1a, assuming an electron to be initially in the Floquet state $\psi_{\mathbf{k}}$. Solving the Schrödinger equation (4) within the first order of the conventional perturbation theory Landau_3 , we arrive at the amplitude of the one-vertex process,
@@
Applying Eq. (Anomalous radiation reaction in a circularly polarized field), one can find the total power that an electron radiates,
@@
which is identical to the classical Larmor formula, $P_{0}=(2e^{2}/3c^{3})\dot{\mathbf{v}}^{2}$, where the electron velocity $\mathbf{v}$ is defined by Eq. (3). As to the recoil force acting on the electron (the momentum transferred to the electron per unit time due to the photon emission),
@@
where $\gamma=1/\sqrt{1-{v}^{2}/c^{2}}$ is the Lorentz factor. Indeed, substituting the velocity (3) into Eq. (15) and averaging it over the field period, $2\pi/\omega$, we arrive at the same Eq. (14). As a consequence, Eq. (14) exactly coincides also with the Landau-Lifshitz equation describing the radiation reaction in an electromagnetic field Landau_2 . To extend the developed theory for an electron in an electromagnetic wave, one needs to make the replacement $\omega t\rightarrow\omega t-\mathbf{q}_{0}\mathbf{r}$ in the vector potential (1), where $\mathbf{q}_{0}=(0,\,0,\,\omega/c)$ is the wavevector of the wave. It can be easily proved that the Floquet functions (2) with the same replacement stay eigenfunctions of the Hamiltonian $\hat{\cal H}_{0}$ in the non-relativistic limit ($v/c,\,\hbar\omega/m_{e}c^{2}\ll 1$). As an only consequence of such a modification of them, the replacement of the delta function, $\delta^{3}(\mathbf{k}-\mathbf{k}^{\prime}\pm\mathbf{q})\rightarrow\delta^{3}(\mathbf{k}-\mathbf{k}^{\prime}\pm\mathbf{q}+m\mathbf{q}_{0})$, should be done in the matrix element (Anomalous radiation reaction in a circularly polarized field). Physically, this means that emission of each photon at the main radiation harmonic with $m=1$ is accompanied by transfer of the momentum $\hbar\mathbf{q}_{0}$ from the wave to the electron (the Compton scattering). Therefore, the photon drag force,
@@
occurs in addition to the recoil force (14), where $\mathbf{n}_{0}=\mathbf{q}_{0}/q_{0}$ is the unit wavevector. As expected, Eq. (16) is identical to the classical radiation pressure force acting on an electron under an electromagnetic wave Landau_2 ; Bohm .
It follows from the aforesaid that the one-vertex emission process pictured in Fig. 1a leads to exactly the same physical results as the classical electrodynamics. To find the first quantum correction to them, one needs to take into account the one-loop process pictured in Fig. 1b. Solving the Schrödinger equation (4) within the third order of the conventional perturbation theory Landau_3 , we arrive at the amplitude of the one-loop process,
@@
is the QED correction to the probability (Anomalous radiation reaction in a circularly polarized field) arisen from the one-loop process SM . Applying Eq. (Anomalous radiation reaction in a circularly polarized field), the total power emitted by an electron can be written as
@@
where $\mathbf{n}=\mathbf{r}/r=(\sin\theta\cos\varphi,\,\sin\theta\sin\varphi,\,\cos\theta)$ is the unit radius vector, $\mathbf{F}_{\|}$ is the normal classical recoil force (14) directed oppositely to the electron velocity $\mathbf{v_{k}}$,
@@
is the anomalous recoil force directed perpendicularly to the velocity $\mathbf{v_{k}}$, and $\mathbf{L}$ is the unit vector along the field angular momentum, which defines the field polarization: The vector $\mathbf{L}=(0,\,0,\,1)$ describes the clockwisely polarized field (1), whereas the vector $\mathbf{L}=(0,\,0,\,-1)$ corresponds to the counterclockwise field polarization.
It should be noted that the emission processes described by the loop and loopless Feynman diagrams make substantially different contributions to the radiation reaction (Anomalous radiation reaction in a circularly polarized field). Mathematically, this follows from the fact that the anomalous recoil force (25) appears due to the singularities of the probability amplitude (Anomalous radiation reaction in a circularly polarized field), which arise from virtual photons SM . In contrast to the loop diagrams, the loopless ones do not involve virtual photons and, therefore, their amplitudes are devoid of the singularities. As a consequence, the two emission processes pictured in Fig. 1 describe the radiation reaction (Anomalous radiation reaction in a circularly polarized field) accurately within the leading order of electron-photon interaction: The one-vertex process makes main contribution to the normal recoil force (14), whereas the one-loop process is responsible for the anomalous force (Anomalous radiation reaction in a circularly polarized field). It should be stressed that the one-vertex photon emission pictured in Fig. 1a is the direct emission process which does not involve intermediate electron-photon states and, therefore, is similar physically to the classical process of electromagnetic emission. As a consequence, it results in the classical recoil force (14), where the Planck constant, $\hbar$, is cancelled. As to the one-loop photon emission pictured in Fig. 1b, it is the indirect process going through the intermediate state with the virtual photon $\mathbf{q}^{\prime}$. It has no analogue within classical electrodynamics and, therefore, yields the anomalous recoil force (25) involving the fine structure constant, $\alpha=e^{2}/\hbar c$, which is of purely quantum nature.
@@
Evolution of the electron velocity $\mathbf{v_{k}}$ under the force (Anomalous radiation reaction in a circularly polarized field) can be described by the Newton equation, $m_{e}\dot{\mathbf{v}}_{\mathbf{k}}=\mathbf{F}_{\|}+\mathbf{F}_{\perp}$, which leads to the electron trajectory pictured schematically in Fig. 3. Since the quantum recoil force, $\mathbf{F}_{\perp}$, is directed perpendicularly to the electron velocity, it does not decelerate an electron in contrast to the classical recoil force, $\mathbf{F}_{\|}$, but bends its trajectory similarly to the Lorentz force acting on an electron in a magnetic field. Such an analogy has the deep physical nature since both a high-frequency circularly polarized field and a stationary magnetic field break the time-reversal symmetry. As a consequence, a circularly polarized field can induce various magnetic-like effects, including the spin polarization Kibis_2022 , bandgap opening in solids Kibis_2010 ; Wang_2013 ; Sie_2015 ; Cavalleri_2020 , persistent currents in conductors Kibis_2011 ; Melnikov_2021 , etc. It should be stressed that the spatial asymmetry of photon emission responsible for the quantum recoil force (25) — which is described by the last term in Eq. (Anomalous radiation reaction in a circularly polarized field) — also arises from the broken time-reversal symmetry. Therefore, the discussed anomalous radiation reaction acting on a charged particle perpendicularly to its velocity takes place in any electromagnetic field with nonzero angular momentum $\mathbf{L}$ (circularly and elliptically polarized fields) and vanishes only in the case of linearly polarized field with $\mathbf{L}=0$. It follows from the aforesaid, particularly, that Eq. (25) should be considered as a quantum correction to the Landau-Lifshitz equation Landau_2 describing the classical radiation reaction force in an electromagnetic field of most general form. It should be noted also that the angular momentum of a circularly polarized field, $\mathbf{L}$, is codirectional to the angular momentum of a charge rotating under the field. This means that the quantum recoil force (25) is similar phenomenologically to the classical Magnus force, $\mathbf{F}_{M}\propto[\mathbf{L}\times\mathbf{v}]$, acting on a body which rotates with the angular momentum $\mathbf{L}$ and makes forward movement with the velocity $\mathbf{v}$ in a medium with friction (see, e.g., Ref. Clancy_book, ). Treating the classical recoil force (14) as a radiation friction, the quantum recoil force (25) can be considered qualitatively as a specific kind of the Magnus force arisen from this friction acting on a rotating charge.
@@
Although the present theory is developed for the homogeneous field (1), the force (25) very weakly depends on spatial profile of the field if its characteristic scale much exceeds the radius of circular electron trajectory, $r_{0}=|v_{0}|/\omega$. As to the time profile of the field (the laser pulse duration), its characteristic scale should much exceed the radiative life time (12). Since the radiative life time, $\tau$, must be large enough to satisfy the condition $\omega\tau\gg 1$, it is most suitable to detect the force (25) with non-relativistic laser fields which can be easily generated in the continuous-wave regime. Discussing dependence of the force (25) on the field strength, $E_{0}$, it should be noted that the Floquet functions (2) are exact solutions of the non-relativistic Schrödinger problem for an electron in the circularly polarized background field (1) of any strength. Therefore, the developed theory is applicable to the background field of the strength $E_{0}$ satisfying only the non-relativistic condition, $v_{0}/c=eE_{0}/m_{e}\omega c\ll 1$. To keep the ratio $v_{0}/c$ to be small enough within applicability of the developed non-relativistic theory, one can consider laser fields of the strength $E_{0}$ up to $10^{11}$ V/m. Estimating the force (25) quantitatively, it should be noted that the electric field amplitude $E_{0}$ can reach $10^{10}$ V/m for an electromagnetic wave generated by an ordinary hundred-kilowatt laser with the wavelengths $\lambda\sim 10^{-6}$ m. In such a field, the quantum force (25) yields the electron acceleration $a_{\perp}=F_{\perp}/m_{e}\sim 10^{3}(v_{\mathbf{k}}/c)$ m/s2, which can be macroscopically large even for non-relativistic electrons satisfying the condition $v_{\mathbf{k}}\ll c$. Thus, the discussed anomalous radiation reaction is the rare case of macroscopic QED effect which can be observable for slow electrons in laser fields of non-relativistic intensity. As to analysis of fast electrons moving in ultra-strong fields, it needs the relativistic extension of the developed Floquet theory based on the Dirac equation SM .
@@
Conclusion—It is shown theoretically that the emission of photons by an electron rotating under a high-frequency circularly polarized electromagnetic field yields the quantum recoil force acting on the electron perpendicularly to the velocity of its forward movement. In contrast to the classical Lorentz-Abraham-Dirac recoil force directed oppositely to the velocity, it does not decelerate the electron but substantially modifies its trajectory. Such an anomalous radiation reaction arises from the one-loop QED correction to the photon emission and can manifest itself in strong laser fields.
@@
Acknowledgments—This work was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” and the Ministry of Science and Higher Education of the Russian Federation (Project FSUN-2023-0006).